# Con4m: Context-aware Consistency Learning Framework for Segmented Time Series Classification

### v1

@@ 4.3 Label Substitution Experiment
Since ambiguous boundaries are inherent to SEEG data and the majority voting procedure is costly, we limit this procedure to only one high-quality testing group in the label disturbance experiment. Besides, on the SEEG data, Con4m modifies approximately $10\%$ of the training labels, which is a significant proportion. Therefore, it is necessary to further evaluate the effectiveness of our label harmonization process on SEEG data. Specifically, we train the TSC baselines based on the harmonized labels generated by Con4m and observe to what extent the performance of TSC models is improved. As shown in Figure 4(a), PatchTST and TimesNet, employing deep learning architectures, are more susceptible to label inconsistency, so they obtain more significant performance improvement ($4.11\%$ and $7.53\%$). Unlike modified PatchTST that considers the contextual data information across consecutive segments, TimesNet only focuses on the independent segments, thus having a more dramatic improvement. In contrast, MiniRocket achieves only a $1.68\%$ increase, indicating that MiniRocket is more robust with a non-deep learning-based simple random feature mapping.
@@ 4.4 Ablation Experiment
Figure 5: Case study for a continuous time interval in SEEG testing set.
@@ 4.5 Case Study
We present a case study to provide a specific example that illustrates how Con4m works for MVD in Figure 5. We show comparative visualization results for the predictions in a continuous time interval in the SEEG testing set. In SEEG data, we assign the label of normal segments as 00 and that of seizures as $1$. As the figure shows, Con4m demonstrates a more coherent narrative by constraining the prediction behavior and aligning with the contextual data information. In contrast, Scale-T, Sel-CL and MiniRocket exhibit noticeably interrupted and inconsistent predictions. MS-TCN2 fails to identify normal segments. More impressively, Con4m accurately identifies the consistent boundary within the time interval spanning across two classes. This verifies that the label consistency framework can harmonize the boundaries more effectively. Refer to Appendix H for more cases.
@@ 5 Conclusion and Discussion
In this work, we focus on the raw time series MVD for segmented time series classification (TSC) tasks, demonstrating unique challenges that are overlooked by existing mainstream TSC models. We first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Based on the insights, we introduce contextual prior knowledge of data locality and label coherence to guide the model to focus on contextual information more conducive to discriminating consecutive segments in segmented TSC tasks. Leveraging effective contextual information, a label consistency learning framework Con4m is proposed to progressively harmonize inconsistent labels during training. Extensive experiments validate the superior performance achieved by Con4m and highlight the effectiveness of the proposed consistent label training framework. Our work still has some limitations. We have solely focused on analyzing and designing end-to-end supervised models. Further exploration of large-scale models would be challenging yet intriguing. When faced with more diverse label behaviors, the function fitting module needs to engage in more selection and design of basis functions. Nevertheless, our work brings new insights to the TSC domain, re-emphasizing the importance of the inherent temporal dependence of time series.
@@ Appendix F Implementation Details of Con4m
The non-linear encoder $g_{\text{enc}}$ used in Con4m is composed of three 1-D convolution layers. The number of kernels vary across different data and you can find corresponding parameters in the default config file of our source code. We construct the Con-Transformer based on the public codes implemented by HuggingFace [^1] https://github.com/huggingface/transformers/blob/v4.25.1/src/transformers/models/bert/modeling_bert.py. We set $d$=$128$ and the dimension of intermediate representations in FFN module as $256$ for all experiments. The number of heads and dropout rate are set as $8$ and $0.1$ respectively. Since we observe that one-layer Con-Attention can fit the data well, we do not stack more layers to avoid overfitting. Note that Con4m consists of two Con-Transformers and we use two Con-Attention layers. The model is optimized using Adam optimizer [34] with a learning rate of $1e$-$3$ and weight decay of $1e$-$4$, and the batch size is set as $64$. We build our model using PyTorch 2.0.0 [48] with CUDA 11.8. And the model is trained on a workstation (Ubuntu system 20.04.5) with 2 CPUs (AMD EPYC 7H12 64-Core Processor) and 8 GPUs (NVIDIA GeForce RTX 3090). You can find more technical details in our source code attached in the supplementary materials.
@@
Appendix G Full Results
The full results of the label disturbance experiment are listed in Table 45 and 6. For fNIRS, we first divide the data into $4$ groups by subjects and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct cross-validation experiments. Therefore, there are $C_{4}^{2}\times C_{2}^{1}=12$ experiments in total. Similarly, we divide the Sleep data into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Therefore, we carry out $C_{3}^{1}\times C_{2}^{1}=6$ experiments. For SEEG data, we follow the same setting as fNIRS. Notice that for SEEG data, inconsistent labels already exist in the raw data. We obtain a high-quality testing group by using a majority voting procedure to determine the boundaries. Then we leave the testing group aside and only change the validation group to report the mean value of $C_{3}^{2}=3$ experiments. All the experimental results are listed in lexicographical order according to the group name composition. We also report the mean value and standard derivation of all experiments. Specifically, we use the STDEVA to estimate standard deviation based on a sample of data.

### v2

@@ 4.3 Label Substitution Experiment
Since ambiguous boundaries are inherent to SEEG data and the majority voting procedure is costly, we limit this procedure to only one high-quality testing group in the label disturbance experiment. Besides, on the SEEG data, Con4m modifies approximately $10\%$ of the training labels, which is a significant proportion. Therefore, it is necessary to further evaluate the effectiveness of our label harmonization process on SEEG data. Specifically, we train the TSC baselines based on the harmonized labels generated by Con4m and observe to what extent the performance of TSC models is improved. As shown in Figure 3(b), PatchTST and TimesNet, employing deep learning architectures, are more susceptible to label inconsistency, so they obtain more significant performance improvement ($4.11\%$ and $7.53\%$). Unlike modified PatchTST that considers the contextual data information across consecutive segments, TimesNet only focuses on the independent segments, thus having a more dramatic improvement. In contrast, MiniRocket achieves only a $1.68\%$ increase, indicating that MiniRocket is more robust with a non-deep learning-based simple random feature mapping.
4.4 Ablation Experiment
@@ 4.5 Case Study
Figure 4: Case study for a continuous time interval in SEEG testing set. The C-score, introduced by the ClaSP model [19], assessing the ability of models to recognize segmentation boundaries by measuring the trade-off between precision (correctly identified change points) and recall (finding all true change points).
We present a case study to provide a specific example that illustrates how Con4m works for MVD in Figure 4. We show comparative visualization results for the predictions in a continuous time interval in the SEEG testing set. In SEEG data, we assign the label of normal segments as 00 and that of seizures as $1$. As the figure shows, Con4m demonstrates a more coherent narrative by constraining the prediction behavior and aligning with the contextual data information. In contrast, Scale-T, Sel-CL and MiniRocket exhibit noticeably interrupted and inconsistent predictions. MS-TCN2 fails to identify normal segments. More impressively, Con4m accurately identifies the consistent boundary within the time interval spanning across two classes. We also utilize the C-score proposed by the ClaSP model [19] to assess the segmentation capability of models. The C-score reflects the ability of models to recognize segmentation boundaries by measuring the trade-off between precision (correctly identified change points) and recall (finding all true change points), ensuring the model captures meaningful transitions without over-segmenting or missing important splits. We compute the scores for each model across three sets of experiments on SEEG data and take the average. Notably, Con4m outperforms the other models significantly, while MS-TCN2, specifically designed for segmentation tasks, also achieves impressive scores. This verifies that the label consistency framework can harmonize the boundaries more effectively. Refer to Appendix H for more cases.
@@ 5 Conclusion and Discussion
In this work, we focus on the raw time series MVD for segmented time series classification (TSC) tasks, demonstrating unique challenges that are overlooked by existing mainstream TSC models. We first formally demonstrate that valuable contextual information enhances the discriminative power of classification instances. Based on the insights, we introduce contextual prior knowledge of data locality and label coherence to guide the model to focus on contextual information more conducive to discriminating consecutive segments in segmented TSC tasks. Leveraging effective contextual information, a label consistency learning framework Con4m is proposed to progressively harmonize inconsistent labels during training. Extensive experiments validate the superior performance achieved by Con4m and highlight the effectiveness of the proposed consistent label training framework. Our work still has some limitations. We have solely focused on analyzing and designing end-to-end supervised models. Further exploration of large-scale models would be challenging yet intriguing. Con4m is a combination of segmentation and classification, both of which are fully supervised. Exploring its application in unsupervised segmentation tasks is worthwhile. When faced with more diverse label behaviors, the function fitting module needs to engage in more selection and design of basis functions. Nevertheless, our work brings new insights to the TSC domain, re-emphasizing the importance of the inherent temporal dependence of time series.
Acknowledgments and Disclosure of Funding
This work was partially supported by National Natural Science Foundation of China (No. 62322606, No. 62441605).
@@ Appendix D Details of Datasets
HHAR. The HHAR (Heterogeneity Human Activity Recognition) dataset [7] captures sensor data from multiple smart devices to explore the impact of device heterogeneity on human activity recognition. It involves measurements from accelerometers and gyroscopes, which are two common motion sensors found in smartphones and smartwatches. The dataset includes six types of human activities: ‘Biking’, ‘Sitting’, ‘Standing’, ‘Walking’, ‘Stairs Up’, and ‘Stairs Down’. For each activity, both 3-axis accelerometer and 3-axis gyroscope readings are recorded, resulting in six key features per time step. The data was collected from five different device types, consisting of four smartwatches and eight smartphones, across nine participants. Each device samples data at its highest possible frequency, generally around $50$Hz. For data preprocessing, the accelerometer and gyroscope signals are aligned based on their timestamps, normalized within each channel, and divided by the devices, following the procedure outlined in Gagnon-Audet et al. [21]. To ensure balanced samples, we combined data from the Galaxy S3 Mini, LG watch, and Gear watch into a single group.
@@ Appendix F Implementation Details of Con4m
The non-linear encoder $g_{\text{enc}}$ used in Con4m is composed of three 1-D convolution layers. The number of kernels vary across different data and you can find corresponding parameters in the default config file of our source code. We construct the Con-Transformer based on the public codes implemented by HuggingFace [^1] https://github.com/huggingface/transformers/blob/v4.25.1/src/transformers/models/bert/modeling_bert.py. We set $d$=$128$ and the dimension of intermediate representations in FFN module as $256$ for all experiments. The number of heads and dropout rate are set as $8$ and $0.1$ respectively. Since we observe that one-layer Con-Attention can fit the data well, we do not stack more layers to avoid overfitting. Note that Con4m consists of two Con-Transformers and we use two Con-Attention layers. The model is optimized using Adam optimizer [38] with a learning rate of $1e$-$3$ and weight decay of $1e$-$4$, and the batch size is set as $64$. We build our model using PyTorch 2.0.0 [52] with CUDA 11.8. And the model is trained on a workstation (Ubuntu system 20.04.5) with 2 CPUs (AMD EPYC 7H12 64-Core Processor) and 8 GPUs (NVIDIA GeForce RTX 3090).
Appendix G Full Results
The full results of the label disturbance experiment are listed in Table 456 and 7. For fNIRS, we first divide the data into $4$ groups by subjects and follow the $2$ training-$1$ validation-$1$ testing ($2$-$1$-$1$) setting to conduct cross-validation experiments. Therefore, there are $C_{4}^{2}\times C_{2}^{1}=12$ experiments in total. Similarly, we divide the HHAR and Sleep data into $3$ groups and follow the $1$-$1$-$1$ experimental setting. Therefore, we carry out $C_{3}^{1}\times C_{2}^{1}=6$ experiments. For SEEG data, we follow the same setting as fNIRS. Notice that for SEEG data, inconsistent labels already exist in the raw data. We obtain a high-quality testing group by using a majority voting procedure to determine the boundaries. Then we leave the testing group aside and only change the validation group to report the mean value of $C_{3}^{2}=3$ experiments. All the experimental results are listed in lexicographical order according to the group name composition. We also report the mean value and standard derivation of all experiments. Specifically, we use the STDEVA to estimate standard deviation based on a sample of data.
@@ Appendix G Full Results
0 1 41.39 34.82 41.92 52.89 45.80 30.88 46.79 52.71 47.69 47.40 45.80 52.37
2 77.19 55.28 53.24 75.76 89.33 52.88 80.21 98.20 93.87 76.48 95.08 95.04
3 43.71 42.61 45.85 54.11 44.75 42.92 46.89 47.60 46.38 48.14 49.43 48.45
4 81.92 82.37 65.32 75.00 92.82 55.04 77.44 96.50 81.12 81.04 93.65 98.78
5 90.67 79.96 70.51 94.37 88.03 70.61 94.05 93.02 87.18 91.70 94.63 94.09
6 83.89 80.07 63.73 69.91 71.70 59.66 68.27 83.61 81.73 67.06 88.05 93.02
Avg 69.79 62.52 56.76 70.34 72.07 52.00 68.94 78.61 73.00 68.64 77.77 80.29
Std 21.56 21.08 11.51 15.47 22.00 13.74 19.01 22.67 20.63 18.01 23.52 23.26
20 1 42.39 44.86 38.64 47.97 45.96 30.52 47.59 51.25 46.59 45.54 48.25 54.82
2 68.54 64.63 54.97 71.37 83.16 39.71 81.73 97.72 90.03 70.52 95.35 93.31
3 46.72 47.95 38.36 49.72 48.64 43.00 51.86 50.40 47.72 53.70 51.06 45.06
4 87.95 77.25 66.51 69.95 85.91 41.34 76.36 91.20 86.85 73.44 93.33 96.37
5 82.97 68.53 69.22 86.79 87.36 68.77 88.61 91.35 89.62 85.33 90.48 90.61
6 71.76 62.32 55.47 54.14 70.13 49.41 64.65 79.70 72.84 67.55 81.81 91.38
Avg 66.72 60.92 53.86 63.32 70.19 45.46 68.47 76.94 72.28 66.02 76.71 78.59
Std 18.63 12.38 13.20 15.26 18.77 12.95 16.55 21.05 20.45 14.29 21.48 22.49
40 1 39.17 42.50 37.82 47.65 44.86 30.87 52.92 48.83 48.60 46.60 50.21 50.78
2 59.62 59.00 51.47 66.27 73.60 42.93 72.53 92.05 94.60 67.78 94.41 91.36
3 54.82 46.54 42.00 48.52 43.99 39.89 52.23 49.13 47.87 55.59 48.69 53.33
4 72.94 78.15 60.46 62.46 87.26 46.10 76.04 75.99 80.98 74.54 93.49 91.57
5 75.12 75.85 57.81 77.54 79.70 62.93 88.42 93.46 86.63 85.48 87.43 82.62
6 72.08 62.57 54.23 53.03 71.13 51.42 63.43 79.79 78.15 64.05 81.56 83.48
Avg 62.29 60.77 50.63 59.25 66.76 45.69 67.60 73.21 72.81 65.67 75.97 75.52
Std 13.94 14.64 8.95 11.68 18.18 10.87 14.13 19.95 19.85 13.74 21.06 18.58
Table 5: Full results of the label disturbance experiment on HHAR data. The best results are in bold and we underline the second best results. The worst results are denoted in italics.
TAS TSC NLL TSC&NLL